# Epicyclic motion of charged particles around a weakly magnetized Kiselev black hole

### v1

@@ Abstract
We investigate the motion of charged particles evolving around a magnetized Kiselev black hole, in the weak magnetic field approximation. The effective potential allows us to study the bound motion and the stable circular orbits. We analyze the impact of combined quintessence and magnetic fields on the epicyclic frequencies. Finally, we examine the periapsis shift and gravitational Larmor precession pointing out differences from the Ernst or Kiselev spacetimes.
@@ 1 Introduction
A central part of this work is devoted to the study of circular and quasi-circular orbits. We derive the conditions for the existence of stable circular motion in the equatorial plane and analyze how the ISCO radius depends on the models parameters. In this respect, we identify parameter ranges for which stable circular orbits exist and provide a systematic classification of allowed angular momenta and energies. Also, we investigate the quasi-harmonic oscillations around stable circular orbits, by computing the radial and latitudinal epicyclic frequencies and exploring the impact of quintessence and magnetic fields on these characteristic frequencies [30]. Particular attention is given to the behavior of the frequencies near the ISCO and near the critical radius beyond which no stable orbits exist. These results are especially relevant in the context of high-frequency quasi-periodic oscillations observed in accreting black hole systems [31]. Finally, we examine relativistic precessional effects, including periapsis precession and gravitational Larmor (nodal) precession pointing out features that have no analogue in Ernst spacetime [8].
@@ 2 Magnetized Kiselev solution
with $\Lambda=1+B_{0}^{2}r^{2}\sin^{2}\theta$, where $B_{0}$ is a constant related to the value of the magnetic induction on the axis of symmetry. In the followings, for simplicity, we shall denote the metric function $f(r)$ by $f$. The magnetic field is oriented orthogonal to the equatorial plane of the black hole being generated by the electromagnetic potential:
$A_{\phi}\,=\frac{B_{0}r^{2}\sin^{2}\theta}{\Lambda}.$ (2)
@@
where $w$ is the equation of state parameter and $k$ is the positive quintessence parameter, which is related to the quintessence energy density. In order to have an accelerated expansion of the Universe, the equation of state parameter $w$ should be in the interval $w\in[-1,-1/3]$.
Thus, the axially-symmetric geometry considered in the present work is sourced by an anisotropic fluid and also by an electromagnetic field sourced by this fluid. In order to understand the effects of the magnetizing technique on the fluid solution see [29]. One may notice that, for $B_{0}=0$, the relation (1) with (3) is the usual Kiselev line element, while for $k=0$, we recover the Ernst solution [7], also known as the Schwarzschild–Melvin solution.
@@ 2.2 Equations of motion in the weak field approximation
The real positive roots of the equation (17) situated between the horizons determine the maxima, minima and saddle points of the effective potential. As expected, in the equatorial plane where the critical points corresponding to the maximum or minimum are located, the left hand side of the equation (18) vanishes. In order to find the solution situated outside the equatorial plane, we impose that the equations (17) and (18) have at least one common solution. Using Cramer’s rule one finds that the off-equatorial critical point is a solution of $f^{\prime}=0$, being given by the relation
@@
These off-equatorial critical points correspond to saddle points, as they cannot represent local extrema of the effective potential. This behaviour has no equivalent in Ernst spacetime, where all critical points are confined to the equatorial plane [7, 9]. In order to verify that the pair $(r_{s},\theta_{s})$ corresponds to a saddle point, we are going to compute the Hessian determinant which reads
$D(r_{s},\theta_{s})=\left.\frac{\partial^{2}V}{\partial r^{2}}\frac{\partial^{2}V}{\partial\theta^{2}}-\frac{\partial^{2}V}{\partial r\partial\theta}\right|_{r=r_{s},\theta=\theta_{s}}=\frac{48Mwb^{2}}{r_{s}}f_{s}\cos^{2}\theta_{s}$
with $f_{s}=f(r=r_{s})$. Since the paramater $w$ is negative, the function $D$ is a negative quantity and therefore $(r_{s},\theta_{s})$ is indeed a saddle point of $V$. The existence of such critical points outside the equatorial plane arises solely due to the presence of quintessence.
@@ 4.2 Illustrative examples
In the followings, let us discuss the dependence of $E_{min}^{2}$ and $E_{s1,s2}^{2}$ in terms of the parameters $k$ and $b$.
@@
One may notice in the figures that the particle’s trajectory can be curly and this happens in cases where $L$ and $b$ have the same sign and $\dot{\phi}$ given by (13) changes sign in $r_{0}$ given by
@@ 5.1 Circular orbits in the equatorial plane
One may notice that the right hand side of the above expression is vanishing in $r=r_{*}=\sqrt{2M/k}$, with $r_{*}$ being the critical point defined in (29), and also in
@@ 1-1 $(3M,6M)$ $0<k<\frac{M}{r^{3}}$
In view of the analysis developed in this section, one may conclude by saying that the radius $r_{ISCO}$ is strongly depending on the model parameters $b$ and $k$ which have to be in their allowed ranges given in (44) and table 3. An illustrative example is shown in figure 9. For $k=0$ and $b=0$, one obtains the well known result $r_{ISCO}=6M$. For the Ernst spacetime (the blue plots), all values of $r_{ISCO}$ are located below $r=6M$ [9]. However, in the presence of quintessence, values of $r_{ISCO}$ bigger than $6M$ are possible, the maximum value of $r_{ISCO}$ being $r_{*}$ given in (29). For a fixed value of $b$, $r_{ISCO}$ is increasing with $k$, while for a given value of $k$, as $|b|$ increases, $r_{ISCO}$ decreases.
@@ 6.1 Fundamental frequencies
One may notice that the latitudinal frequency does not depend on $b$ and this can be understood by the fact that the Lorentz force is acting only on the radial direction. However, in the presence of quintessence, $\Omega_{\theta}^{2}$ decreases with $k$. On the other hand, the radial frequency $\Omega_{r}$ is characterizing the oscillatory motion in the equatorial plane under the influence of the Lorentz force, attracting gravitation and repulsive quintessence. Thus, $\Omega_{r}^{2}$ has a more complicated form, being dependent on both $b$ and $k$. For $b=0$ and $k=0$, the expressions corresponding to the Schwarzschild case are recovered, i.e.
@@ 6.1.1 The positive magnetic parameter
Even though the expressions of $\omega_{\theta}^{2}$ and $\omega_{\phi}$ are same same as for the Ernst spacetime [9], the contribution of quintessence is encoded in the angular momentum (33). As it can be noticed in the figures 11 and 12, both $\omega_{\theta}$ and $\omega_{\phi}$ are monotonically decreasing with $r$ and this behavior is more prominent as $k$ or $b$ increases. For $k\neq 0$, there is a critical value of $r$, denoted by $r_{*}$ and defined in (29), where $\omega_{\theta}$ vanishes and the frequency $\omega_{\phi}$ changes the sign. However, the value $r_{*}$ is inside the range where $\omega_{r}$ is defined and, in view of the analysis developed in the previous section, the allowed range of $r$ allowing the existence of stable circular orbits is $r\in[r_{ISCO},r_{*}]$. In what it concerns $r_{ISCO}$ where $\omega_{r}$ is vanishing, the conditions given in table 3 should be satisfied.
@@ 6.1.2 The negative magnetic parameter
Secondly, let us mention that the same analysis can be done for $b<0$. Since the magnetic field and angular momentum have opposite signs, the Lorentz force is attracting the charged particle towards the OzOz-axis. In the right panels of the figures 13, 14 and 15, one may see that all frequencies are increasing as $|b|$ increases.
@@
In the presence of quintessence, similarly to the previous case, all frequencies are decreasing as $k$ increases (see the left panels of the figures 13, 14, 15). For $k=0$, both $\omega_{r}$ and $\omega_{\phi}$ are approaching the Larmor frequency $\omega_{L}$, the first one from below and the other from above. However, by comparing the plots in the left panels of the figures 10 and 13, we notice that, for $b<0$, the values of $\omega_{r}$ and the ranges of $r$ are much smaller since the charged particle is attracted towards the black hole by both gravitation and Lorentz force.
@@
Also, by comparing the plots in the right panels of the figures 11 and 15, we see that, for $b<0$, the frequency $\omega_{\phi}$ is always positive and its values are significantly increased.
@@ 6.2.1 Periapsis shift
The same analysis can be done for negative values of $b$. For a comparison, in the the right panel of the figure 16, we have changed the sign of $b$ but kept the same values of the other parameters. As a result of the attractive Lorentz force, the plot of $\omega_{\phi}$ have moved upwards and the one of $\omega_{r}$ downwards. Since the plot of $\omega_{\phi}$ is above the one of $\omega_{r}$, it means that all orbits with $r\in[r_{ISCO},r_{*})$ have a prograde precession. Once $k$ is decreasing, the two curves come closer to each other, approaching $\omega_{L}$, one from below and the other from above (see the left panel of figures 13 and 15). Even though $\Delta\phi$ is decreasing for $r$ moving from $r_{ISCO}$ to $r_{*}$, the periastron shift remains positive for all the trajectories.
@@ 6.2.2 The gravitational Larmor precession
If the orbit is slightly inclined to the equatorial plane, a gravitational Lorentz precession may appear []. This is characterized by the difference between $\omega_{\phi}$ and $\omega_{\theta}$ as
@@
Conclussions of the discussion on the signs of $\Delta\phi$ and $\Delta\theta$ presented in this section are given in table 4, for positive and negative values of $b$ and $r$ in its allowed range $r\in[r_{ISCO},r_{p}]$.
@@ 6.3 Illustrative examples of quasi-circular orbits
As a first example of a perturbed circular orbit of radius $r_{c}$, let us consider the stable trajectory with $r_{c}$ close to $r_{*}=\sqrt{2M/k}$ represented in the figure 18. In the three-dimensional representation given in the left panel, one may notice the quasi-harmonic oscillations both around $r_{c}$ and with respect to the equatorial plane. In the projection on the equatorial plane (a quarter is given in the right panel), the region of stable circular orbits is between the cyan and black circles which correspond to $r_{ISCO}$ and $r_{*}$ respectively. The values of $L$ and $E^{2}$ are in the shaded regions represented in the figures 7 and 8. The unperturbed circular orbit of radius $r_{c}$ is represented by the red circle which is very close to the green circle of radius $r_{0}=\sqrt{L/b}$ where $\phi$ is changing the sign. As expected, because $r_{c}>r_{p}$, the orbit has a negative periastron shift and is curly since one of the turning points is beyond $r_{0}$.
@@
In view of the analysis developed in section 5, let us mention that circular orbits with $r_{c}>r_{*}$ are unstable. Such an example is given in the figure 19 where one may also notice that the orbit of the particle escaping towards the cosmological horizon is curled outward the black hole. This happens because the repulsive quintessence contribution becomes dominant for $r>r_{*}$. On the other hand, in figure 18, we see an opposite behaviour for the particle’s stable orbit with radius $r<r_{*}$ which is curled toward the black hole as a result of gravity dominance.
Figure 19: A 3D representation of an unstable quasi-circular trajectory with $r_{c}>r_{*}$ and its projection on the $xOy$ plane. The numerical values are: $M=1$, w=2/3w=-2/3, $k=0.015$, $b=0.09$, $L=20$, $E^{2}=0.65$. The dark red circle is the horizon $r_{-}=2.064$, the cyan circle is the ISCO{ISCO} with $r_{ISCO}=4.22$, the black circle corresponds to $r_{*}=11.547$, the green circle of radius $r_{0}=14.90$ is separating the region of curly trajectories from the non-curly one and the red circle is the unperturbed circular orbit with $r_{c}=15.06$.

### v2

@@ Abstract
We investigate the motion of charged particles evolving around a magnetized Kiselev black hole, in the weak magnetic field approximation. The effective potential allows us to study the bound motion and the stable circular orbits. We analyse the impact of combined quintessence and magnetic fields on the epicyclic frequencies. Finally, we examine the periapsis shift and gravitational Larmor precession pointing out differences from the Ernst or Kiselev spacetimes.
@@ 1 Introduction
A central part of this work is devoted to the study of circular and quasi-circular orbits. We derive the conditions for the existence of stable circular motion in the equatorial plane and analyse how the ISCO radius depends on the model’s parameters. In this respect, we identify parameters ranges for which stable circular orbits exist and provide a systematic classification of allowed angular momenta and energies. Also, we investigate the quasi-harmonic oscillations around stable circular orbits, by computing the radial and latitudinal epicyclic frequencies and exploring the impact of quintessence and magnetic fields on these characteristic frequencies [30]. Particular attention is given to the behavior of the frequencies near the ISCO and near the critical radius beyond which no stable orbits exist. These results are especially relevant in the context of high-frequency quasi-periodic oscillations observed in accreting black hole systems [35]. Finally, we examine relativistic precessional effects, including periapsis precession and gravitational Larmor (nodal) precession pointing out features that have no analogue in Ernst spacetime [8].
@@ 2 Magnetized Kiselev solution
with $\Lambda=1+B_{0}^{2}r^{2}\sin^{2}\theta$, where $B_{0}$ is a constant related to the value of the magnetic induction on the axis of symmetry. In the followings, for simplicity, we shall denote the metric function $f(r)$ by $f$. The electromagnetic potential
$A_{\phi}\,=\frac{B_{0}r^{2}\sin^{2}\theta}{\Lambda}$ (2)
is generating a magnetic field which is axially aligned, breaking the spherical symmetry of the spacetime. The components of the magnetic field are given by the expressions
$B_{r}=-\frac{4B_{0}r\sin\theta\cos\theta}{\Lambda},\>B_{\phi}=\frac{4B_{0}r^{2}\sin^{2}\theta}{\Lambda}.$ (3)
@@
where $w$ is the equation of state parameter. In the above relation, $k$ is the quintessence parameter, which is related to the quintessence fluid energy density as:
$\rho^{0}=-\frac{3kw}{8\pi r^{3(w+1)}}$ (5)
while the components of the anisotropic pressures are: $p_{r}^{0}=-\rho^{0}$ and
$p_{t}^{0}\equiv p_{\theta}^{0}=p_{\phi}^{0}=-\frac{3(3w+1)kw}{16\pi r^{3(w+1)}}.$ (6)
In the presence of the magnetic field, the fluid energy density and pressures are modified and their explicit expressions can be found in [29]. However, for a weak magnetic field, which is the case considered in the present work, the expressions of energy density and pressures can be approximated to the ones given in (5) and (6). In what it concerns the values of the quintessence parameter $k$, some constraints have been obtained by comparing the theoretical results with data provided by the EHT Collaboration for M87* and SgrA* [31]. For example, in [32], the authors are assuming that these objects can be modelled as Kerr supermassive black holes surrounded by quintessence. From the shadows observables, they found that the upper limits on the quintessence parameter are: $kM=0.0358$ for M87* BH and $kM=0.011$ for SgrA* BH. The latter result agrees with the recent estimation for a Frolov BH surrounded by quintessence $kM\leq 0.012$ [21].
In what it concerns the equation of state parameter $w$, this should belong to the interval $w\in[-1,-1/3]$ in order to induce an accelerated expansion of the Universe. The particular value w=2/3w=-2/3 has been intensively used in theoretical models since, in this case, one has a linear contribution in the metric function and the whole analysis becomes simpler leading to significant analytical results. However, a special attention should be also given to the value w=1w=-1 which is favoured by cosmological data.
Finally, an important parameter which has a strong influence on the particle’s trajectory is the magnetic induction $B_{0}$. As we are going to see, the model’s parameters $B_{0}$ and $k$ compete against each other and one has to find ranges of these parameters for which the effective potential has a minimum value between two maxima and particles with suitable energies can be trapped in bound orbits.
One may notice that, for $B_{0}=0$, the relation (1) with (4) is the usual Kiselev line element, while for $k=0$, we recover the Ernst solution [7], also known as the Schwarzschild–Melvin solution.
@@ 2.2 Equations of motion in the weak field approximation
The real positive roots of the equation (20) situated between the horizons determine the maxima, minima and saddle points of the effective potential. As expected, in the equatorial plane where the critical points corresponding to the maximum or minimum are located, the left hand side of the equation (21) vanishes. In order to find the solutions situated outside the equatorial plane, we impose that the equations (20) and (21) have at least one common solution. Using Cramer’s rule one finds that the off-equatorial critical point is a solution of $f^{\prime}=0$, being given by the relation
@@
These off-equatorial critical points correspond to saddle points, as they cannot represent local extrema of the effective potential. This behaviour has no equivalent in Ernst spacetime, where all critical points are confined to the equatorial plane [7, 9].
In order to verify that the pair $(r_{s},\theta_{s})$ which is a critical point of the differentiable real function $V(r,\theta)$ corresponds to a saddle point, one has to compute the Hessian matrix associated to the potential $V$ . The corresponding determinant reads
$D(r_{s},\theta_{s})=\left.\frac{\partial^{2}V}{\partial r^{2}}\frac{\partial^{2}V}{\partial\theta^{2}}-\left(\frac{\partial^{2}V}{\partial r\partial\theta}\right)^{2}\right|_{r=r_{s},\theta=\theta_{s}}=\frac{48Mwb^{2}}{r_{s}}f_{s}\cos^{2}\theta_{s}$ (25)
with $f_{s}=f(r=r_{s})$. Since the parameter $w$ is negative and $f_{s}>0$, one may conclude by saying that the function $D(r_{s},\theta_{s})$ is a negative quantity and therefore $(r_{s},\theta_{s})$ is indeed a saddle point of $V$ [33]. The existence of such critical points outside the equatorial plane arises solely due to the presence of quintessence.
@@ 4.1 Conditions for bound orbits to exist
Let us underline that $r_{*}$ can be named as the static radius (as in [34]) where the gravitational attraction is compensated by the quintessence repulsion. This value of the radial coordinate will play an important role in separating the region corresponding to $r<r_{*}$ where one may have stable orbits, from the one with $r>r_{*}$ where the particle’s trajectories are always unstable. In our analysis, we shall consider $kM\ll 1$ so that the cosmological horizon is far from the black hole’s horizon and from the region where the particle’s bound trajectories are confined. This assumption agrees with other theoretical estimates on the upper limits of the quintessence parameter [21]. As an example, let us consider the special case corresponding to w=2/3w=-2/3 for which the two horizons have the analytical expressions given in (8). For $kM\ll 1$, the expressions of the horizons can be approximated to
$r_{-}\approx 2M\;,\quad r_{+}\approx\frac{1}{k}$
and one has the relation $r_{-}<r_{*}=\sqrt{2M/k}\ll r_{+}$.
@@ 4.2 Illustrative examples
In the followings, let us discuss the dependence of $E_{min}^{2}$ and $E_{s1,s2}^{2}$, defined in the previous subsection, on the parameters $k$ and $b$.
@@
If the particle starts its motion from a saddle point, its trajectory is either a capture one, as in the figure 5, or an escape one, as in the figure 6.
Figure 5: Left panel. Plot of the effective potential (18) and the capture particle trajectory in the $(x,z)$- plane. The particle with the energy $E^{2}=0.60$ starts its motion from the saddle point $(r_{s},\theta_{s})=(10,0.68)$. Right panel. 3D plot of the capture trajectory of a particle moving in the potential represented in the left panel. The gray sphere represents the horizon $r_{-}=2.09$. The values of the parameters are: $M=1$, w=2/3w=-2/3, $k=0.02$, $b=0.10$ and $L=4$.
Figure 6: Left panel. Plot of the effective potential (18) and the escape particle trajectory in the $(x,z)$- plane. The particle starts its motion from the saddle point $(r_{s},\theta_{s})=(8.17,0.89)$. The energy $E^{2}=0.51$ is represented by the light blue horizontal plane. Right panel. 3D plot of the escape trajectory of the particle moving in the potential represented in the left panel. The gray sphere represents the horizon $r_{-}=2.14$. The values of the parameters are: $M=1$, w=2/3w=-2/3, $k=0.03$, $b=0.10$ and $L=4$.
Finally, one may notice in the figures that the particle’s trajectory can be curly and this happens in cases where $L$ and $b$ have the same sign and $\dot{\phi}$ given by (16) changes sign in $r_{0}$ given by
@@ 5.1 Circular orbits in the equatorial plane
One may notice that the right hand side of the above expression is vanishing in $r=r_{*}=\sqrt{2M/k}$, with $r_{*}$ being static radius defined in (33), and also in
@@ 1-1 $(3M,6M)$ $0<k<\frac{M}{r^{3}}$
In view of the analysis developed in this section, one may conclude by saying that the stable circular trajectories, if exist, are in the region $r\in[r_{ISCO},r_{*}]$, where $r_{*}$ is depending only on the quintessence parameters $w$ and $k$ (see the formula (33)). In what it concerns the radius $r_{ISCO}$, this is strongly depending on both parameters $b$ and $k$ which have to be in their allowed ranges given in (48) and table 3. An illustrative example is shown in figure 11. For $k=0$ and $b=0$, one obtains the well known result $r_{ISCO}=6M$. For the Ernst spacetime (the blue plots), all values of $r_{ISCO}$ are located below $r=6M$ [9]. However, in the presence of quintessence, values of $r_{ISCO}$ bigger than $6M$ are possible, the maximum value of $r_{ISCO}$ being $r_{*}$ given in (33). For a fixed value of $b$, $r_{ISCO}$ is increasing with $k$, while for a given value of $k$, as $|b|$ increases, $r_{ISCO}$ decreases.
@@ 6.1 Fundamental frequencies
One may notice that the latitudinal frequency does not depend on $b$ and this can be understood by the fact that the Lorentz force is acting only on the radial direction. However, in the presence of quintessence, $\Omega_{\theta}^{2}$ decreases with $k$.
At $r=r_{*}=\sqrt{2M/k}$, the quantity $\Omega_{\theta}$ is vanishing. Thus, one may conclude by saying that the static radius $r=r_{*}$ is separating the region $r<r_{*}$ where $\Omega_{\theta}^{2}>0$ and the circular trajectories are stable under fluctuations in the transverse direction from the one corresponding to $r>r_{*}$ where $\Omega_{\theta}^{2}<0$. In this case, there are no vertically stable orbits and the particle is accelerated along the $Oz$-axis toward infinity (see the equations (51)).
On the other hand, the radial frequency $\Omega_{r}$ is characterizing the oscillatory motion in the equatorial plane under the influence of the Lorentz force, attracting gravitation and repulsive quintessence. Thus, $\Omega_{r}^{2}$ has a more complicated form, being dependent on both $b$ and $k$. For $b=0$ and $k=0$, the expressions corresponding to the Schwarzschild case are recovered, i.e.
@@ 6.1.1 The positive magnetic parameter
Even though the expressions of $\omega_{\theta}^{2}$ and $\omega_{\phi}$ are same same as for the Ernst spacetime [9], the contribution of quintessence is encoded in the angular momentum (37). As it can be noticed in the figures 13 and 14, both $\omega_{\theta}$ and $\omega_{\phi}$ are monotonically decreasing with $r$ and this behavior is more prominent as $k$ or $b$ increases. For $k\neq 0$, there is a critical value of $r$, denoted by $r_{*}$ and defined in (33), where $\omega_{\theta}$ vanishes and the frequency $\omega_{\phi}$ changes the sign. However, the value $r_{*}$ is inside the range where $\omega_{r}$ is defined and, in view of the analysis developed in the previous section, the allowed range of $r$ allowing the existence of stable circular orbits is $r\in[r_{ISCO},r_{*}]$. In what it concerns $r_{ISCO}$ where $\omega_{r}$ is vanishing, the conditions given in table 3 should be satisfied.
@@ 6.1.2 The negative magnetic parameter
Secondly, let us mention that the same analysis can be done for $b<0$. Since the magnetic field and angular momentum have opposite signs, the Lorentz force is attracting the charged particle towards the OzOz-axis. In the right panels of the figures 15, 16 and 17, one may see that all frequencies are increasing as $|b|$ increases.
@@
In the presence of quintessence, similarly to the previous case, all frequencies are decreasing as $k$ increases (see the left panels of the figures 15, 16, 17). For $k=0$, both $\omega_{r}$ and $\omega_{\phi}$ are approaching the Larmor frequency $\omega_{L}$, the first one from below and the other from above. However, by comparing the plots in the left panels of the figures 12 and 15, we notice that, for $b<0$, the values of $\omega_{r}$ and the ranges of $r$ are much smaller since the charged particle is attracted towards the black hole by both gravitation and Lorentz force.
@@
Also, by comparing the plots in the right panels of the figures 13 and 17, we see that, for $b<0$, the frequency $\omega_{\phi}$ is always positive and its values are significantly increased.
@@ 6.2.1 Periapsis shift
The same analysis can be done for negative values of $b$. For a comparison, in the the right panel of the figure 18, we have changed the sign of $b$ but kept the same values of the other parameters. As a result of the attractive Lorentz force, the plot of $\omega_{\phi}$ have moved upwards and the one of $\omega_{r}$ downwards. Since the plot of $\omega_{\phi}$ is above the one of $\omega_{r}$, it means that all orbits with $r\in[r_{ISCO},r_{*})$ have a prograde precession. Once $k$ is decreasing, the two curves come closer to each other, approaching $\omega_{L}$, one from below and the other from above (see the left panel of figures 15 and 17). Even though $\Delta\phi$ is decreasing for $r$ moving from $r_{ISCO}$ to $r_{*}$, the periastron shift remains positive for all the trajectories.
@@ 6.2.2 The gravitational Larmor precession
If the orbit is slightly inclined to the equatorial plane, a gravitational Larmor precession may appear [38]. This is characterized by the difference between $\omega_{\phi}$ and $\omega_{\theta}$ as
@@
Conclusions of the discussion on the signs of $\Delta\phi$ and $\Delta\theta$ presented in this section are given in table 4, for positive and negative values of $b$ and $r$ in its allowed range $r\in[r_{ISCO},r_{p}]$.
@@ 6.3 Illustrative examples of quasi-circular orbits
As a first example of a perturbed circular orbit of radius $r_{c}$, let us consider the stable trajectory with $r_{c}$ close to $r_{*}=\sqrt{2M/k}$ represented in the figure 20. In the three-dimensional representation given in the left panel, one may notice the quasi-harmonic oscillations both around $r_{c}$ and with respect to the equatorial plane. In the projection on the equatorial plane (a quarter is given in the right panel), the region of stable circular orbits is between the cyan and black circles which correspond to $r_{ISCO}$ and $r_{*}$ respectively. The values of $L$ and $E^{2}$ are in the shaded regions represented in the figures 9 and 10. The unperturbed circular orbit of radius $r_{c}$ is represented by the red circle which is very close to the green circle of radius $r_{0}=\sqrt{L/b}$ where $\phi$ is changing the sign. As expected, because $r_{c}>r_{p}$, the orbit has a negative periastron shift and is curly since one of the turning points is beyond $r_{0}$.
@@
In view of the analysis developed in section 5, let us mention that circular orbits with $r_{c}>r_{*}$ are unstable. Such an example is given in the figure 21 where one may also notice that the orbit of the particle escaping towards the cosmological horizon along the $z$-axis is curled toward the black hole. This happens because the repulsive quintessence contribution becomes dominant for $r>r_{*}$. On the other hand, in figure 20, we see an opposite behaviour for the particle’s stable orbit with radius $r<r_{*}$ which is curled outward the black hole as a result of gravity dominance.
Figure 21: A 3D representation of an vertically unstable quasi-circular trajectory with $r_{c}>r_{*}$ and its projection on the $xOy$ plane. The numerical values are: $M=1$, w=2/3w=-2/3, $k=0.015$, $b=0.09$, $L=20$, $E^{2}=0.65$. The dark red circle is the horizon $r_{-}=2.064$, the cyan circle is the ISCO{ISCO} with $r_{ISCO}=4.22$, the black circle corresponds to $r_{*}=11.547$, the green circle of radius $r_{0}=14.90$ is separating the region of curly trajectories from the non-curly one and the red circle is the unperturbed circular orbit with $r_{c}=15.06$.